OpenStudy (anonymous):

I have this problem: b-4/6=b/2 I am confused on how to solve it.

jimthompson5910 (jim_thompson5910):

Is the equation \[\Large \frac{b-4}{6} = \frac{b}{2}\] ???

OpenStudy (anonymous):

Yes that is it.

jimthompson5910 (jim_thompson5910):

Start off by cross multiplying \[\Large \frac{b-4}{6} = \frac{b}{2}\] \[\Large 2(b-4) = 6b\]

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (anonymous):

I do not understand.

jimthompson5910 (jim_thompson5910):

If you have a/b = c/d, you cross multiply to get a*d = b*c

jimthompson5910 (jim_thompson5910):

example: 1/2 = 3/6 1*6 = 2*3

OpenStudy (anonymous):

Thank you!!!!

jimthompson5910 (jim_thompson5910):

yw

jimthompson5910 (jim_thompson5910):

keep going to solve for b and tell me what you get

OpenStudy (anonymous):

2(b-4)=6b - 2b2b-8=6b-2b -8=4b 4b-8=4b4b b=-2

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thank You!

jimthompson5910 (jim_thompson5910):

you're welcome